OpenStudy (anonymous):

Simplify by removing the factors of 1. 405v^5y^10/75v^3y^8. show steps I am having trouble with this one still.

OpenStudy (anonymous):

405v^5y^10/75v^3y^8 you can reduce 405/75 by dividing both numerator and denominator by 15 that is 405/75=27/5, for the powers of v just subtract them like v^5/v^3=v^2 and for the powers of y also subtract them,like y^10/y^8=y^2 and therefore 405v^5y^10/75v^3y^8 =(27/5)v^2y^2 ans.... or =(27v^2y^2)/5 ans....